### v2

# Embeddable partial groups

Abstract.
We record a folklore theorem that says a partial group embeds in a group if and only if each word has at most one possible multiplication, regardless of choice of parenthesization. We further investigate the partial groups which are exemplars of non-embeddability. Finally we show that a partial groupoid embeds in a groupoid if and only if its reduction embeds in a group.
This work was supported by a grant from the Simons Foundation (#850849, PH) and by a grant from the Simons Foundation International (SFI-MPS-TSM-00014188, JL). PH was partially supported Louisiana Board of Regents through the Board of Regents Support fund LEQSF(2024-27)-RD-A-31. ES is grateful for travel support from TU Nachwuchsring at the RPTU Kaiserslautern-Landau.
A partial group (in Chermak’s sense [5]) is embeddable if it is an impartial subgroup of some group. There is an evident obstruction to embeddability: if there is a word $w$ in the elements of the partial group and two total parenthesizations of that word such that the successive binary multiplications specified by the two parenthesizations exist and are different, then it is impossible to embed in any group. We show here that the converse holds, a result essentially going back to Baer (Remark˜4). This theorem has been spoken about since the early days of partial groups. For example, it was a topic of discussion in July 2011 at the masterclass “Fusion systems and $p$-local group theory” in Copenhagen.
The corresponding result without assuming invertibility of elements is false, but the theorem is also true in the many-object case (partial groupoids), and it is no more effort to work at this level of generality. Our proof uses well-known material about the reflection $\tau\colon\mathsf{sSet}\to\mathsf{Cat}$ of simplicial sets into categories going back to Gabriel & Zisman [8].
We also construct a collection of universal counterexamples to embeddability, one for each suitable pair of parenthesizations of a generic length $n$ word. Using these, the category of embeddable partial groupoids may be concretely described as an orthogonality class in the category of partial groupoids. We compute the degree, in the sense of [12], of these universal counterexamples.
This paper uses the formalism of symmetric (simplicial) sets, an enhancement of simplicial sets which is suitable for modeling invertibility. Symmetric sets have proved to be a convenient home for working with partial groups [13]. Let $\Upsilon$ be the category whose objects are the sets $n]=\{0,1,\dots,n\}$ for $n\geq 0$, with arbitrary functions as morphisms. A symmetric set is just a functor $\Upsilon^{\textup{op}}\to\mathsf{Set}$. By composing with the inclusion $\Delta\to\Upsilon$, we obtain a functor $\mathsf{Sym}\to\mathsf{sSet}$ which takes a symmetric set to its underlying simplicial set.
1. Embeddability
The nerve functor $N\colon\mathsf{Cat}\to\mathsf{sSet}$ admits a left adjoint $\tau$ which sends a simplicial set $X$ to the category $\tau X$ which is formed by first taking the free category on the underlying reflexive directed graph and then imposing the relation $(d_{0}\sigma)\circ(d_{2}\sigma)=d_{1}\sigma$ for every $\sigma\in X_{2}$. In particular, $\tau X$ only depends on the 2-skeleton of $X$.
The fundamental groupoid functor is the composite
$\mathsf{sSet}$$\mathsf{Cat}$$\mathsf{Gpd},$$\scriptstyle{\tau}\scriptstyle{\bot}$
where the left adjoint $\mathsf{Cat}\to\mathsf{Gpd}$ inverts all morphisms. The fundamental group at a vertex $x\in X_{0}$ is the set of endomorphisms of $x$ in the fundamental groupoid. The composite $\mathsf{Sym}\to\mathsf{sSet}\to\mathsf{Cat}$ already lands in the category of groupoids, so the fundamental groupoid of a symmetric set $X$ is just $\tau X$.
If $X$ is a simplicial set, $\mathbf{W}(X)$ is the graded set of compatible words in $X_{1}$. That is, $\mathbf{W}(X)_{0}=X_{0}$ and
$\mathbf{W}(X)_{n}=X_{1}\times_{X_{0}}X_{1}\times_{X_{0}}\dots\times_{X_{0}}X_{1}=\{x_{0}\xrightarrow{f_{1}}x_{1}\xrightarrow{f_{2}}\cdots\xrightarrow{f_{n}}x_{n}\}.$
We say $X$ is edgy if the Segal map $\mathscr{E}_{n}\colon X_{n}\to\mathbf{W}(X)_{n}$ picking out principal edges is injective for all $n\geq 1$ [13, Definition 1.7]. An edgy simplicial set is a kind of partial category, with partially-defined $n$-ary composition given by the span $\mathbf{W}(X)_{n}\hookleftarrow X_{n}\to X_{1}$, whose rightward leg is the composition of inner face maps. A symmetric set is called spiny if its underlying simplicial set is edgy [13, §3]; we also call spiny symmetric sets partial groupoids. Each groupoid is a partial groupoid via the fully faithful nerve functor $N\colon\mathsf{Gpd}\to\mathsf{Sym}$ [11, 4.1], whose left adjoint is $\tau$. In this paper a partial group is a reduced spiny symmetric set [13, Theorem A].
From now on we are only interested in edgy simplicial sets. Write $\mathbf{W}^{+}(X)\subseteq\mathbf{W}(X)$ for the set of positive-length words. We define a transitive relation $\rightsquigarrow$ on $\mathbf{W}^{+}(X)$ using the partially-defined inner face maps [12, §4.1]. Namely, if $w=(f_{1},\dots,f_{n})$ and there is a 2-simplex $f_{i}|f_{i+1}]\in X_{2}$, we write
$w\rightsquigarrow(f_{1},\dots,f_{i+1}\circ f_{i},\dots,f_{n})$
where $f_{i+1}\circ f_{i}=d_{1}[f_{i}|f_{i+1}]$. The relation $\rightsquigarrow$ is the transitive closure of this operation. In particular, if $w\rightsquigarrow w^{\prime}$ then on lengths we have $|w|>|w^{\prime}|$. The set of morphisms of $\tau X$ is the quotient of $\mathbf{W}^{+}(X)$ by the equivalence relation generated by $\rightsquigarrow$. Write $w]\subseteq\mathbf{W}^{+}(X)$ for the equivalence class of $w\in\mathbf{W}^{+}(X)$.
Suppose $f,g\in X_{1}$ become equal in $\tau X$. There is a zig-zag
[…]
between $f$ and $g$. In general, we don’t have any control over the length of this zig-zag between $f$ and $g$.
Example 1.
Consider the 2-dimensional simplicial set with eight nondegenerate 2-simplices and thirteen nondegenerate 1-simplices, as displayed below.
$1{1^{\prime}}2{2^{\prime}}1{1^{\prime}}2{2^{\prime}}0{0}3{3}0{0}3{3}$$\scriptstyle{b}$$\scriptstyle{c}$$\scriptstyle{b}$$\scriptstyle{e}$$\scriptstyle{c}$$\scriptstyle{a}$$\scriptstyle{f}$$\scriptstyle{d}$$\scriptstyle{a}$$\scriptstyle{h}$$\scriptstyle{q}$$\scriptstyle{r}$$\scriptstyle{q}$$\scriptstyle{t}$$\scriptstyle{r}$$\scriptstyle{p}$$\scriptstyle{h}$$\scriptstyle{s}$$\scriptstyle{p}$$\scriptstyle{g}$
This simplicial set is edgy. We have a zigzag $f\mathrel{\reflectbox{$\rightsquigarrow$}}(a,b,c)\rightsquigarrow h\mathrel{\reflectbox{$\rightsquigarrow$}}(p,q,r)\rightsquigarrow g$ but no shorter one will do to relate $f$ and $g$. We can vary this construction to require a zigzag of arbitrarily high length.
Theorem 2.
Let $X$ be a partial groupoid and $f,g\in X_{1}$. If $[f]=[g]$ in $\tau X$, then there is a word $w\in\mathbf{W}^{+}(X)$ with $f\mathrel{\reflectbox{$\rightsquigarrow$}}w\rightsquigarrow g$.
Proof.
Suppose $n\geq 1$ and we have a zigzag of elements of $\mathbf{W}^{+}(X)$ as follows.
[…]
Set
$w=\begin{cases}w_{1}w_{3}^{-1}w_{5}w_{7}^{-1}\cdots w_{2n-3}^{-1}w_{2n-1}&\text{if $n$ is odd}\\ w_{1}w_{3}^{-1}w_{5}w_{7}^{-1}\cdots w_{2n-1}^{-1}w_{2n}&\text{if $n$ is even}\end{cases}$
and consider the ‘odd reduction’ that replaces all $w_{2k-1}w_{2k+1}^{-1}$ with $w_{2k}w_{2k}^{-1}$ for $1\leq k<n$ odd, and the ‘even reduction’ that replaces all $w_{2k-1}^{-1}w_{2k+1}$ with $w_{2k}^{-1}w_{2k}$ for $2\leq k<n$ even. Suppose $n$ is odd. Then the odd reduction of $w$ takes the form
$w_{1}w_{3}^{-1}w_{5}w_{7}^{-1}\cdots w_{2n-3}^{-1}w_{2n-1}\rightsquigarrow w_{2}w_{2}^{-1}w_{6}w_{6}^{-1}\cdots w_{2n-4}^{-1}w_{2n-1}$
which further reduces to $w_{2n-1}$. The even reduction of $w$ takes the form
$w_{1}w_{3}^{-1}w_{5}w_{7}^{-1}\cdots w_{2n-3}^{-1}w_{2n-1}\rightsquigarrow w_{1}w_{4}^{-1}w_{4}w_{8}^{-1}\cdots w_{2n-2}^{-1}w_{2n-2}$
which further reduces to $w_{1}$. Thus when $n$ is odd we have
$w_{0}\mathrel{\reflectbox{$\rightsquigarrow$}}w_{1}\mathrel{\reflectbox{$\rightsquigarrow$}}w\rightsquigarrow w_{2n-1}\rightsquigarrow w_{2n}.$
Next suppose $n$ is even. The odd reduction of $w$ takes the form
$w_{1}w_{3}^{-1}w_{5}w_{7}^{-1}\cdots w_{2n-3}w_{2n-1}^{-1}w_{2n}\rightsquigarrow w_{2}w_{2}^{-1}w_{6}w_{6}^{-1}\cdots w_{2n-2}w_{2n-2}^{-1}w_{2n}$
which further reduces to $w_{2n}$. The even reduction of $w$ takes the form
$w_{1}w_{3}^{-1}w_{5}w_{7}^{-1}\cdots w_{2n-3}w_{2n-1}^{-1}w_{2n}\rightsquigarrow w_{1}w_{4}^{-1}w_{4}w_{8}^{-1}\cdots w_{2n-4}^{-1}w_{2n-1}^{-1}w_{2n}$
which further reduces to $w_{1}w_{2n-1}^{-1}w_{2n}\rightsquigarrow w_{1}w_{2n}^{-1}w_{2n}\rightsquigarrow w_{1}$. Thus when $n$ is even we have
$w_{0}\mathrel{\reflectbox{$\rightsquigarrow$}}w_{1}\mathrel{\reflectbox{$\rightsquigarrow$}}w\rightsquigarrow w_{2n}.$
If $[f]=[g]$ then we have a zig-zag as in (1) with $w_{0}=f$ and $w_{2n}=g$ (noting that when $f=g\colon a\to b$ we may use the word $w_{1}=(\operatorname{id}_{a},f)$). We then have $f\mathrel{\reflectbox{$\rightsquigarrow$}}w\rightsquigarrow g$. ∎
Recall the partially defined inner face maps $d_{i}\colon\mathbf{W}(X)_{n}\nrightarrow\mathbf{W}(X)_{n-1}$ for $0<i<n$. A word $w\in\mathbf{W}(X)_{n}$ is mean if there are sequences $i_{1},\dots,i_{n-1}$ and $j_{1},\dots,j_{n-1}$ with $0<i_{k},j_{k}<k+1$ such that $d_{i_{1}}\dots d_{i_{n-1}}(w)$ and $d_{j_{1}}\dots d_{j_{n-1}}(w)$ are defined and are different, and kind otherwise. (For example, $d_{1}d_{2}d_{3}(w)\neq d_{1}d_{2}d_{2}(w)$.) We call each one-simplex arising from iterated inner faces of a mean word sad. All other one-simplices are called happy.
Corollary 3.
Let $X$ be a partial groupoid. The following are equivalent:
  1. (1)
    $X$ embeds in a groupoid.
  2. (2)
    $X$ embeds in its fundamental groupoid via the unit map $X\to N\tau X$.
  3. (3)
    All one-simplices are happy.
  4. (4)
    All words are kind.
Proof.
If $X$ embeds in a groupoid, i.e. if $X\subseteq NG$ for some groupoid $G$, then the inclusion is equal to $X\to N\tau(X)\to N\tau(NG)=NG$, so $X$ embeds in its fundamental groupoid. Spininess implies that embeddability of $X$ in $N\tau X$ can be detected at the level of 1-simplices. By the previous theorem, $f\in X_{1}$ is happy if and only if $f]\cap X_{1}=\{f\}$. Thus $X$ embeds in its fundamental groupoid if and only if all 1-simplices are happy. ∎
Remark 4.
Theorem˜2 and Corollary˜3 are many-object versions of Corollary 1 and Corollary 2 of [3, §4]. Indeed, a 2-skeletal partial group [15] is essentially the same thing as what Baer termed an ‘add which is self-reflexive in the strict sense’: in the forward direction one sends $X$ to $X_{1}$ equipped with its partial binary multiplication, while in the reverse direction one builds a partial group whose set of 2-simplices is the set of multipliable pairs (for details see [14]). Baer’s theorems were later rediscovered by Tamari [22, §III], who called these partial magmas monoïdes symétriques. The one-object version of Theorem˜2 was later known as Tamari’s One Mountain Theorem, e.g. [4, Theorem 2.a].
Definition 5.
A partial groupoid $X$ is embeddable if $X\to N\tau X$ is a monomorphism.
For a partial group $X$ this definition agrees with that from the first sentence of the article, as an impartial subgroup of $X$ is precisely a nonempty symmetric subset.
Theorem 6.
The category of embeddable partial groupoids is a reflective subcategory of the category of partial groupoids. (Similarly for partial groups.)
Proof.
Use the epi-mono factorization of the unit natural transformation $\eta\colon\operatorname{id}\Rightarrow N\tau$, i.e. the reflection is given as the first map $X\twoheadrightarrow\overline{X}\rightarrowtail N\tau X$. ∎
Example 7.
Consider the 2-dimensional simplicial set with eight nondegenerate 2-simplices and thirteen nondegenerate edges, as depicted below.
[…]
This simplicial set is edgy. Notice
$f\mathrel{\reflectbox{$\rightsquigarrow$}}(a,b,c)\rightsquigarrow(a,p)\mathrel{\reflectbox{$\rightsquigarrow$}}(a,m,c^{\prime})\rightsquigarrow(q,c^{\prime})\mathrel{\reflectbox{$\rightsquigarrow$}}(a^{\prime},b^{\prime},c^{\prime})\rightsquigarrow g,$
so $f$ and $g$ represent the same morphism in $\tau X$, but all words are kind. Indeed, $f$ and $g$ are the only parallel edges, hence the only potentially sad edges, but none of the four relations $f\mathrel{\reflectbox{$\rightsquigarrow$}}(a,m,c^{\prime})\rightsquigarrow g$, $(a,b,c)\rightsquigarrow g$, $(a^{\prime},b^{\prime},c^{\prime})\rightsquigarrow f$ hold. Thus the simplicial analogue of Theorem˜2 is false.
Example 8.
The nerve of a pregroup (in the sense of Stallings) is embeddable [19, 21]. Pregroups, as well as known generalizations (e.g. [16, 18]), satisfy an associativity axiom stating that if $ab$ and $bc$ are defined, then $(ab)c$ or $a(bc)$ defined implies $(ab)c=a(bc)$. This axiom need not hold in an embeddable partial group – for instance it does not hold for the reduction and symmetrization of the horn $\Lambda^{3}_{1}\subset\Delta^{3}$.
Example 9.
The free partial group on one generator embeds in $B\mathbb{Z}/3\mathbb{Z}$, but its fundamental group is $\mathbb{Z}$. Thus $\pi_{1}(X)\to G$ need not be injective when $X$ embeds in $G$.
Remark 10.
Fundamental groups of linking systems of saturated fusion systems are of interest in $p$-local group theory [2, §III.7 (8)]. Linking systems are special kinds of transporter systems [20, §3], which are equivalent to special partial groups called localities [5], [9, §2]. The locality $L$ associated with a transporter system $\mathcal{T}$ is a quotient $N\mathcal{T}\twoheadrightarrow L$ of the nerve, and this quotient map is a weak equivalence [10, Appendix]. Theorem˜2 may thus be interpreted in light of the computations of fundamental groups of linking systems that have already been carried out. As one example, $\pi_{1}(L^{c}_{\mathrm{Sol}}(q))=1$ for the centric linking locality of a Benson–Solomon fusion system [6]. By Theorem˜2 any two elements of $L^{c}_{\mathrm{Sol}}(q)$ have some word associating to each.
2. Orthogonality set
In [12] a nonembeddable partial groupoid is constructed, based on the two possible triangulations of the square. This partial groupoid is universal with respect to the property $(fg)h\neq f(gh)$ among the binary compositions in the partial groupoid. The analysis of the previous section relied on more complicated parenthesizations of words, and in general we will need more complicated variations of this example to detect nonembeddability. Recall that (full) parenthesizations of $n$ symbols may equally well be described as a triangulation of an (n+1)(n{+}1)-gon. These are the elements of the Tamari lattice, or the vertices in a Stasheff associahedron.
Figure 1. Flip graphs showing different triangulations. Image by TheMathCat at English Wikipedia, licensed under $C$ BY 4.0. https://en.wikipedia.org/wiki/File:Flip_graphs.svg
We now consider triangulations $T$ of a regular (n+1)(n{+}1)-gon having vertex set $\{0,1,\dots,n\}$. Each such $T$ determines a symmetric subset
$H(T)\subseteq\operatorname{sk}_{2}\Upsilon^{n}\subseteq\Upsilon^{n}$
consisting of precisely those triangles in $T$ (here $\Upsilon^{n}$ denotes the representable functor $\hom_{\Upsilon}(-,[n])\colon\Upsilon^{\textup{op}}\to\mathsf{Set}$). Notice that $\textup{Sp}^{n}\subseteq\Upsilon^{n}$, consisting of those functions $m]\to[n]$ whose image lands in $\{i,i+1\}$ for some integer $i$, is a symmetric subset of $H(T)$ for every such triangulation. We regard $T$ as a set of triples $(i,j,k)$ with $0\leq i<j<k\leq n$, where $i,j,k$ are the vertices appearing in a triangle.
Proposition 11.
Let $T$ and $T^{\prime}$ be two triangulations of a regular (n+1)(n{+}1)-gon. The symmetric set $H(T)\amalg_{\textup{Sp}^{n}}H(T^{\prime})$ is spiny if and only if for each $i=1,\dots,n-1$,
$(i-1,i,i+1)\notin T\cap T^{\prime}.$
Proof.
Let $X=H(T)\amalg_{\textup{Sp}^{n}}H(T^{\prime})$. If $(i-1,i,i+1)\in T\cap T^{\prime}$, then there are distinct 2-simplices having spine $i-1\to i\to i+1$, one coming from $H(T)$ and the other coming from $H(T^{\prime})$, so the pushout of symmetric sets is not spiny.
Suppose $(i-1,i,i+1)\notin T\cap T^{\prime}$ for any $i$. To prove that the 2-dimensional symmetric set $X$ is spiny, it suffices to show that $\mathscr{E}_{2}\colon X_{2}\to X_{1}\times_{X_{0}}X_{1}$ is injective [15, Theorem 9]. As the subobjects $H(T)$, $H(T^{\prime})$ of $X$ are spiny, it suffices to consider $x\in H(T)_{2}$ and $x^{\prime}\in H(T^{\prime})_{2}$ such that $\mathscr{E}_{2}(x)=\mathscr{E}_{2}(x^{\prime})$ in $X$. Since the intersection of $H(T)$ and $H(T^{\prime})$ in $X$ is $\textup{Sp}^{n}$, both edges
$k_{0}$$k_{1}$$k_{2},$$\scriptstyle{f}$$\scriptstyle{g}$
appearing in $\mathscr{E}_{2}(x)$ are in $\textup{Sp}^{n}$, hence $|k_{1}-k_{0}|\leq 1$ and $|k_{2}-k_{1}|\leq 1$. We thus have $\{k_{0},k_{1},k_{2}\}\subseteq\{i-1,i,i+1\}$ for some $i$; by our assumption, equality does not hold. Further, $\{k_{0},k_{1},k_{2}\}\neq\{i-1,i+1\}$ since it would imply either $f$ or $g$ not in $\textup{Sp}^{n}$. We conclude that $\{k_{0},k_{1},k_{2}\}$ is a subset of either $\{i,i-1\}$ or $\{i,i+1\}$, so $x,x^{\prime}\in\textup{Sp}^{n}$. Thus $x=x^{\prime}$ since $\textup{Sp}^{n}$ is spiny. ∎
Definition 12.
If $T$ and $T^{\prime}$ are two triangulations of the regular (n+1)(n{+}1)-gon such that $H(T)\amalg_{\textup{Sp}^{n}}H(T^{\prime})$ is spiny, we write $\operatorname{NA}^{T,T^{\prime}}$ for this partial groupoid. We say that the pair $T,T^{\prime}$ is compatible in this case.
See Figures 2 and 3. Note that $\operatorname{NA}$ from [12, §2.3] is $\operatorname{NA}^{T,T^{\prime}}$ for $T,T^{\prime}$ the two triangulations of a square.
0123401234
Figure 2. Incompatible triangulations of the pentagon
012340123401234
Figure 3. Compatible triangulations of the pentagon and $\operatorname{NA}^{T,T^{\prime}}$
Proposition 13.
A partial groupoid $X$ is embeddable if and only if for each compatible pair $T,T^{\prime}$ of triangulations of an (n+1)(n{+}1)-gon, any map $\operatorname{NA}^{T,T^{\prime}}\to X$ identifies the two long edges $0\to n$.
Proof.
Suppose a map $\operatorname{NA}^{T,T^{\prime}}\to X$ sends the long edges to $f,g\in X_{1}$. The image of $\textup{Sp}^{n}\subset\operatorname{NA}^{T,T^{\prime}}$ produces a word $w\in\mathbf{W}(X)_{n}$ such that $f\mathrel{\reflectbox{$\rightsquigarrow$}}w\rightsquigarrow g$, so $f]=[g]\in\tau X$. If $X$ is embeddable, then $f=g$.
Conversely, suppose $f\neq g\in X_{1}$ with $[f]=[g]$. By Theorem˜2 there is a word $w$ with $f\mathrel{\reflectbox{$\rightsquigarrow$}}w\rightsquigarrow g$. Assume $w$ has the minimal length (necessarily greater than two) with respect to this property. The parenthesizations of $w=u_{1}\cdots u_{k}$ yielding $f$ and $g$ cannot both contain $(u_{i}u_{i+1})$ for any $i=1,\dots,k-1$, or there would exist a word of shorter length. Letting $T$ and $T^{\prime}$ be the triangulations associated to the two parenthesizations, we thus have $T,T^{\prime}$ is a compatible pair by Proposition˜11. By design, we have a map $\operatorname{NA}^{T,T^{\prime}}\to X$ which takes the spine to $w$ and the two long edges to $f,g$. ∎
Let $\textup{cSp}^{n}\subseteq\operatorname{sk}_{1}\Upsilon^{n}$ be the usual spine together with the long edge $0\to n$ (the circular spine). A slight variation on the proof of Proposition˜11 gives the following.
Proposition 14.
Let $T$ and $T^{\prime}$ be two triangulations of a regular (n+1)(n{+}1)-gon. The symmetric set $H(T)\amalg_{\textup{cSp}^{n}}H(T^{\prime})$ is spiny if and only $H(T)\amalg_{\textup{Sp}^{n}}H(T^{\prime})$ is spiny and
$(n-1,n,0),(n,0,1)\notin T\cap T^{\prime}.$
Definition 15.
If $T$ and $T^{\prime}$ are two triangulations of the regular (n+1)(n{+}1)-gon such that $H(T)\amalg_{\textup{cSp}^{n}}H(T^{\prime})$ is spiny, we say the pair $T,T^{\prime}$ is well-behaved and write $\operatorname{A}^{T,T^{\prime}}$ for this partial groupoid.
Notice that if $n>3$ and $T,T^{\prime}$ are adjacent in the flip graph (Figure˜1), then necessarily the pair is either not compatible or not well-behaved.
Lemma 16.
Let $X$ be a partial groupoid, $g,g^{\prime},h,h^{\prime},f\in X_{1}$. Suppose $(f,g)\rightsquigarrow h$ and $(f,g^{\prime})\rightsquigarrow h^{\prime}$. Then $g=g^{\prime}$ if and only if $h=h^{\prime}$. (Similarly if $(g,f)\rightsquigarrow h$ and $(g^{\prime},f)\rightsquigarrow h^{\prime}$.)
Proof.
The forward direction is clear. For the reverse, the assumption implies $[f|g]$ and $f|g^{\prime}]$ are 2-simplices, hence $f^{-1}|f|g]$ and $f^{-1}|f|g^{\prime}]$ are 3-simplices. But then we have
$g=d_{1}d_{1}[f^{-1}|f|g]=d_{1}d_{2}[f^{-1}|f|g]=d_{1}[f^{-1}|h]$
and similarly $g^{\prime}=d_{1}[f^{-1}|h^{\prime}]$, so $h=h^{\prime}$ implies $g=g^{\prime}$. ∎
If $C$ is a category, then an object $x\in C$ is orthogonal to a map $f\colon a\to b$ if for each map $g\colon a\to x$ there is a unique map $h\colon b\to x$ with $g=hf$. The following is a refinement of Theorem˜6 by [1, Theorem 1.39].
Proposition 17.
A partial groupoid $X$ is embeddable if and only if it is orthogonal to the set of maps
$\left\{\operatorname{NA}^{T,T^{\prime}}\to\operatorname{A}^{T,T^{\prime}}\mid T,T^{\prime}\text{ well-behaved}\right\}.$
Proof.
Suppose $T,T^{\prime}$ is a compatible pair of triangulations of an (n+1)(n{+}1)-gon (in particular $n>2$), and suppose we have a map $\operatorname{NA}^{T,T^{\prime}}\to X$ with $T,T^{\prime}$ not well-behaved (we may then conclude $n>3$). We either have $(n-1,n,0)\in T\cap T^{\prime}$ or $(n,0,1)\in T\cap T^{\prime}$. Deleting the vertex $n$ in the first case, or the vertex 0 in the second, yields a compatible pair of triangulations $S,S^{\prime}$ of an $n$-gon, along with the composite map
$\operatorname{NA}^{S,S^{\prime}}\to\operatorname{NA}^{T,T^{\prime}}\to X.$
By Lemma˜16, $\operatorname{NA}^{S,S^{\prime}}\to X$ identifies long edges if and only if $\operatorname{NA}^{T,T^{\prime}}\to X$ identifies long edges. We may continue this process until we arrive at a well-behaved pair. Thus the statement of Proposition˜13 holds with ‘compatible pair’ replaced by ‘well-behaved pair.’ But for a well-behaved pair $T,T^{\prime}$, $\operatorname{NA}^{T,T^{\prime}}\to X$ identifies the long edges if and only if it factors through $\operatorname{A}^{T,T^{\prime}}$. ∎
3. Degree of $\operatorname{NA}^{T,T^{\prime}}$
The degree is an invariant of symmetric sets introduced in [12], which is based on the notion of higher Segal space [7]. As an example, $\deg(X)=1$ if and only if $X$ is a groupoid. In [12, Example 9.8] it was shown that $\deg(\operatorname{NA}^{T,T^{\prime}})=3$ for $T,T^{\prime}$ the two triangulations of the square.
Let $T,T^{\prime}$ be a compatible pair of triangulations of the regular (n+1)(n{+}1)-gon. A cone from $T$ to $T^{\prime}$ is a pair of triangles (i1,i,k)(i-1,i,k) and $(i,i+1,k)$ in $T$ such that (i1,i,i+1)(i-1,i,i+1) is a triangle in $T^{\prime}$. We say that a compatible pair $(T,T^{\prime})$ has a cone if there is a cone from $T$ to $T^{\prime}$, or vice-versa.
Proposition 18.
If a compatible pair of triangulations $(T,T^{\prime})$ does not have a cone, then $\deg(\operatorname{NA}^{T,T^{\prime}})=2$. Otherwise, $\deg(\operatorname{NA}^{T,T^{\prime}})=3$.
Proof.
Let $T,T^{\prime}$ be a compatible pair of triangulations and $X=\operatorname{NA}^{T,T^{\prime}}$. Since $X$ is 2-dimensional, $\deg(X)\leq 3$ by [12, Theorem 9.6]. As $X$ is not embeddable in a groupoid, it cannot be a groupoid, so $1<\deg(X)$. We write $\mathbf{S}(X)_{n}$ for the set of tuples $(f_{1},\dots,f_{n})\in\prod X_{1}$ such that each $f_{i}$ has the same source; since $X$ is edgy there is an injection $\mathscr{B}_{n}\colon X_{n}\hookrightarrow\mathbf{S}(X)_{n}$ (see [12, §2.2]). By [12, Proposition 4.10], $\mathscr{B}_{n}$ is surjective for every $n\geq k+1$ if and only if $\deg(X)\leq k$. We show that $X$ has a cone if and only if $\deg(X)=3$.
If $(T,T^{\prime})$ has a cone then there is a starry word $x=(f_{1},f_{2},f_{3})\in\mathbf{S}(X)_{3}$ as pictured such that $d_{1}x,d_{2}x,d_{3}x\in X_{2}$.
[…]
This starry word $x$ is not in $X_{3}$, since if it were then it would be nondegenerate [12, Lemma 2.13], and $X$ is 2-dimensional. Thus $X$ is not lower $3$-Segal, hence $\deg(X)>2$.
Conversely, suppose that $X$ is not lower $3$-Segal, i.e. $\deg(X)\geq 3$. If we have a starry word $x=(f_{1},\dots,f_{n})$ with $f_{i}=\operatorname{id}_{a}$, then $x\in X_{n}$ if and only if $d_{i}x=(f_{1},\dots,\hat{f}_{i},\dots,f_{n})\in X_{n-1}$. Indeed, if $\sigma\colon[n]\to[n-1]$ is the zero-preserving map defined by
$\sigma(t)=\begin{cases}0&t=0,i\\ t&1\leq t\leq i-1\\ t-1&i+1\leq t\leq n\end{cases}$
then $\sigma^{*}d_{i}x=x$ when $f_{i}=\operatorname{id}_{a}$. A similar argument shows that if $f_{i}=f_{j}$, then $d_{i}x\in X_{n-1}$ if and only if $x\in X_{n}$.
Since $X$ is not lower $3$-Segal, we can choose $n$ minimal such that there exists $x=(f_{1},\dots,f_{n})\in\mathbf{S}(X)_{n}$ with $d_{1}x,d_{2}x,d_{3}x\in X_{n-1}$, but $x\notin X_{n}$ [12, Proposition 4.10]. By the above remarks, no $f_{i}$ is an identity, and $f_{1},f_{2}$, and $f_{3}$ each appear exactly once among the $f_{i}$.
[…]
The starry words $(f_{1},f_{2})$, $(f_{1},f_{3})$, and $(f_{2},f_{3})$ are in $X_{2}$. We thus have the triangles $(a,i_{1},i_{2})$, $(a,i_{1},i_{3})$, and $(a,i_{2},i_{3})$ in $T\cup T^{\prime}$. They cannot all be in $T$ (resp. $T^{\prime}$) since $T$ is a triangulation. Without loss of generality, suppose $(f_{1},f_{2}),(f_{2},f_{3})\in H(T)$ and $(f_{1},f_{3})\in H(T^{\prime})$. Then $f_{1}$ and $f_{3}$ are in $H(T)\cap H(T^{\prime})=\textup{Sp}^{n}$, so we have $i_{1}=a\pm 1$ and $i_{3}=a\mp 1$. We’ve thus constructed a cone. ∎
Below we give an alternative characterization of degree in terms of the ternary associativity condition mentioned in Example˜8: a partial groupoid $X$ satisfies $\diamondsuit$ if its underlying partial binary composition has the property that whenever $ab$ and $bc$ are defined, then $(ab)c$ is defined if and only if $a(bc)$ is defined, with equality when both hold.
Lemma 19.
If a partial groupoid $X$ does not satisfy $\diamondsuit$, then $\deg(X)>2$.
Proof.
If $X$ does not satisfy $\diamondsuit$, then there is a word $w=(c,b,a)\in\mathbf{W}(X)_{3}$ such that $ab$, bcbc, and $a(bc)$ exist, but either $(ab)c$ does not exist or it does and $(ab)c\neq a(bc)$. The faces $d_{0}w=[b|a]$, $d_{1}w=[bc|a]$, and $d_{3}w=[c|b]$ are in $X_{2}$, but $w\notin X_{3}$. Thus $X$ is not lower 3-Segal by [12, Remark 3.17]. ∎
If $G$ is a group with at least four elements, then the 2-skeleton of $BG$ satisfies $\diamondsuit$ and has degree 3, so the converse of the lemma does not always hold.
Proposition 20.
Let $(T,T^{\prime})$ be a compatible pair of triangulations of a regular $(n+1)$-gon and $X=\operatorname{NA}^{T,T^{\prime}}$. Then $X$ satisfies $\diamondsuit$ if and only if $\deg(X)=2$.
Proof.
The backwards direction follows immediately from Lemma˜19. We use Proposition˜18 for the forward direction: if $X$ has degree 3, then $(T,T^{\prime})$ has a cone, and without loss of generality we assume there are triangles (i1,i,k),(i,i+1,k)(i-1,i,k),(i,i+1,k) in $T$ and (i1,i,i+1)(i-1,i,i+1) in $T^{\prime}$. There are edges $c\colon i-1\to i$ and $b\colon i\to i+1$ in $\textup{Sp}^{n}$, $a\colon i+1\to k$, $ab\colon i\to k$, and $(ab)c\colon i-1\to k$ in $H(T)$, and $bc\colon i-1\to i+1$ in $H(T^{\prime})$. If $a(bc)$ does not exist, then $\diamondsuit$ does not hold. Suppose $a(bc)$ exists. As $bc$ is not in $\textup{Sp}^{n}$, hence not in $H(T)$, it must be the case that $a\in\textup{Sp}^{n}$. Since $k\neq i,i+1$, we have $k=i+2$ and (i1,i+1,i+2)(i-1,i+1,i+2) is a triangle of $T^{\prime}$. But in this case $a(bc)\colon i-1\to i+2$ in $H(T^{\prime})$ is necessarily different from $(ab)c\colon i-1\to i+2$ in $H(T)$, so again $\diamondsuit$ fails. ∎
4. Reduction and embeddability
The reduction of a simplicial or symmetric set identifies all of the vertices; if $X$ is a partial groupoid, then its reduction is a partial group [13, Proposition 5.3]. Theorem˜2 has the following consequence. Later we establish the converse.
Lemma 21.
Let $X$ be a partial groupoid. If the reduction of $X$ is an embeddable partial group, then $X$ is embeddable.
Proof.
Given a word $w\in\mathbf{W}^{+}(X)_{n}$, if $f\mathrel{\reflectbox{$\rightsquigarrow$}}w\rightsquigarrow g$ then the source of $f$ and $g$ are equal. In particular, if they are both identities, then $f=g$. If $X$ is not embeddable, then there exists a mean word $w$. By the preceding sentence, the reduction of $w$ is still mean. Thus the reduction $\mathcal{R}X$ of $X$ is not embeddable. ∎
Suppose $C$ is a category and define a monoid $M=M(C)$ whose elements are strings $(f_{1},\dots,f_{n})$ for $n\geq 0$ where
  1. (1)
    $f_{i}\colon a_{i-1}\to b_{i}$ is a nonidentity morphism of $C$, and
  2. (2)
    $a_{i}\neq b_{i}$ for each $i$.
As usual, a string satisfying (1) is reduced, and we call a string satisfying condition (2) jagged.
If we have an arbitrary word which is not reduced, we may delete any identities to make it reduced. Likewise, given a non-jagged word there is a serration process which uses composition to arrive at a jagged word. Of course the reduction of a jagged word may not be jagged, and the serration of a reduced word may not be reduced. Letting $\operatorname{\mathsf{rs}}$ be this iterated reduction-serration procedure, the multiplication of two elements $x=(f_{1},\dots,f_{n})$ and $y=(g_{1},\dots,g_{m})$ is defined to be
$x\bullet y\coloneqq\operatorname{\mathsf{rs}}(yx)=\operatorname{\mathsf{rs}}(g_{1},\dots,g_{m},f_{1},\dots,f_{n}).$
For convenience, we rephrase the definition of this monoid in terms of rewrite systems and confluence [17, §2.1–2.2]. Set $S$ be the set of words in $C_{1}$, and consider the following rewrite rule $\rightarrow$ on $S$:
$(f_{1},\dots,f_{n})$ $\rightarrow(f_{1},\dots,f_{i+1}\circ f_{i},\dots,f_{n})$
$(f_{1},\dots,f_{i-1},\operatorname{id}_{a},f_{i},\dots,f_{n})$ $\rightarrow(f_{1},\dots,f_{i-1},f_{i},\dots,f_{n})$
[…]
Lemma 22.
Each element of $(S,\rightarrow)$ has a unique normal form.
Proof.
The rewriting system $(S,\rightarrow)$ is terminating, since $S$ is graded by length and $\rightarrow$ is strictly decreasing on length. We show below that $(S,\rightarrow)$ is locally confluent, i.e. $x\leftarrow w\rightarrow y$ implies there exists $w^{\prime}$ such that $x\overset{*}{\rightarrow}w^{\prime}\overset{*}{\leftarrow}y$. Newman’s lemma says that a locally confluent, terminating rewrite system is confluent, i.e. $x\overset{*}{\leftarrow}w\overset{*}{\rightarrow}y$ implies there exists $w^{\prime}$ such that $x\overset{*}{\rightarrow}w^{\prime}\overset{*}{\leftarrow}y$. In particular, each element of $S$ has a unique normal form.
To prove that $(S,\rightarrow)$ is locally confluent, it suffices to show given $x\leftarrow w\rightarrow y$ that $x=y$ or there exists $z$ with $x\rightarrow z\leftarrow y$. Write $w=(f_{1},\dots,f_{n})$ and assume $x\neq y$. As there are two types of $\rightarrow$, we have three possibilities to consider. First, suppose $x=(f_{1},\dots,\hat{f}_{i},\dots,f_{n})$ and $y=(f_{1},\dots,\hat{f}_{j},\dots,f_{n})$ with $i<j$ and $f_{i},f_{j}$ are identities. Then
$x\rightarrow(f_{1},\dots,\hat{f}_{i},\dots,\hat{f}_{j},\dots,f_{n})\leftarrow y.$
Next suppose $x=(f_{1},\dots,\hat{f}_{i},\dots,f_{n})$ and $y=(f_{1},\dots,f_{j+1}\circ f_{j},\dots,f_{n})$ where $f_{i}$ is an identity. We cannot have $i=j,j+1$, or else we’d have $x=y$. Setting
$z=\begin{cases}(f_{1},\dots,\hat{f}_{i},\dots,f_{j+1}\circ f_{j},\dots,f_{n})&i<j\\ (f_{1},\dots,f_{j+1}\circ f_{j},\dots,\hat{f}_{i},\dots,f_{n})&i>j+1\end{cases}$
we have $x\rightarrow z\leftarrow y$. Finally, suppose $i<j$ and $x=(f_{1},\dots,f_{i+1}\circ f_{i},\dots,f_{n})$ and $y=(f_{1},\dots,f_{j+1}\circ f_{j},\dots,f_{n})$. Setting
$z=\begin{cases}(f_{1},\dots,f_{i+1}\circ f_{i},\dots,f_{j+1}\circ f_{j},\dots,f_{n})&i\leq j-2\\ (f_{1},\dots,f_{i+2}\circ f_{i+1}\circ f_{i},\dots,f_{n})&i=j-1\end{cases}$
we have $x\rightarrow z\leftarrow y$. ∎
Lemma 23.
Multiplication in $M$ is associative.
Proof.
Given elements $x,y,z\in M$, we have
$\begin{gathered}xyz\overset{*}{\rightarrow}\operatorname{\mathsf{rs}}(xy)z\overset{*}{\rightarrow}\operatorname{\mathsf{rs}}(\operatorname{\mathsf{rs}}(xy)z)\\ xyz\overset{*}{\rightarrow}x\operatorname{\mathsf{rs}}(yz)\overset{*}{\rightarrow}\operatorname{\mathsf{rs}}(x\operatorname{\mathsf{rs}}(yz))\end{gathered}$
and so the terms on the right are normal forms of xyzxyz. By the previous lemma,
$z\bullet(y\bullet x)=\operatorname{\mathsf{rs}}(\operatorname{\mathsf{rs}}(xy)z)=\operatorname{\mathsf{rs}}(x\operatorname{\mathsf{rs}}(yz))=(z\bullet y)\bullet x.\qed$
Proposition 24.
There is a functor $C\to M$ sending a nonidentity element $f\colon a\to b$ to $(f)$ and $\operatorname{id}_{a}\colon a\to a$ to $()$.
Proof.
The functor may uniformly be described by $f\mapsto\operatorname{\mathsf{rs}}(f)$. Given a composable pair $f,g$ of morphisms in $C$, we have
$(g)\bullet(f)=\operatorname{\mathsf{rs}}(f,g)=\operatorname{\mathsf{rs}}(g\circ f).\qed$
Lemma 25.
If $C$ is a category, then $\mathcal{R}C$ embeds in the monoid $M=M(C)$.
Proof.
The functor $C\to M$ factors as $C\to\mathcal{R}C\to M$. The image of $C_{1}$ in $M$ is precisely the set of words of length at most 1, and $(\mathcal{R}C)_{1}\to M$ is injective. ∎
Notice that if $C$ is a groupoid, then $M$ is a group: $(f_{1},\dots,f_{n})^{-1}=(f_{n}^{-1},\dots,f_{1}^{-1})$.
Theorem 26.
A partial groupoid $X$ is embeddable if and only if its reduction is so.
Proof.
Suppose $X$ is embeddable. By factoring the composite $X\to N\tau X\to\mathcal{R}N\tau X$ through $\mathcal{R}X$ and applying Lemma˜25 to the category $\tau X$, we have the following diagram
${N\tau X}X{\mathcal{R}X}$$\mathcal{R}N\tau X$$M(\tau X)$}
The reduction of an injective map is injective, so $\mathcal{R}X$ embeds in the group $M(\tau X)$. The converse is Lemma˜21. ∎
Acknowledgments
We thank Andy Chermak and Rémi Molinier for discussions related to this paper.